User name Calarco
Log entry time 16:02:51 on November 3,2000
Entry number 51416
Shift crew: Calarco, Averett, Punjabi
Beam into hall at about 0800 .... unstable at present
0900 Beam appears to be stable. Moeller measurement crew has arrived.
1000 Moeller measurements continue; 20 minutes of beam during past hour
1100 Moeller measurements continue; only 10 minutes of beam during past hour
1200 Moeller measurements continue; 20 minutes of beam during past hour
1300 Moeller measurements continue; no beam for past hour
Trips coming from injector according to MCC. They are investigating.
1500 Still no beam; have runs 1043 taken with a little bit of beam earlier, 1044
is pure cosmics; 1045 is pure cosmics.
1505 Ended run 1045; MCC to send us some pulsed beam
1600 This shift ends with a wimper; Moeller measurement is only half done